In a race, Car A starts 1 mile behind Car B. Car A is traveling at 55 mph, while Car B is traveling at 40 mph. How long will it take for Car A to overtake Car B?
Found 2 solutions by stanbon, checkley71:
Answer by stanbon(48551)

In a race, Car A starts 1 mile behind Car B. Car A is traveling at 55 mph, while Car B is traveling at 40 mph. How long will it take for Car A to overtake Car B?
-----------
Car A DATA:
Rate = 55 mph ; Distance : x+1 mile ; Time = d/r = (x+1)/55 hrs.
-------------
Car B DATA:
Rate = 40 mph ; Distance : x miles ; Time = d/r = x/40
----------------------------------------
A time = B time
(x+1)/55 = x/40
40(x+1) = 55x
15x = 40
x = 8/3 miles
-------------
Time = x/40 = (8/3)/40
Time = 1/15 hr. = 15 minutes
========================
Cheers,
Stan H.

Answer by checkley71(8403)

THE DIFFERENCE IN SPEEDS IS 55-40=15 MPH
THE DIFFERENCE IN THE MILEAGE IS 1.
AT 15 MPH IT TAKES 1/15 HOUR TO COVER 1 MILE.
1/15 HOUR=X/60 MINUTES
1/15=X/60
15X=60
X=60/15
X=4 MINUTES FOR CAR A TO CATCH CAR B.
40(4/60)=160/60=16/6=8/3=2.667 MILES
55(4/60)=220/60=11/3=3.667 MILES
SO IN 4 SECONDS THE FASTER CAR COVERS 3.667 WHILE THE SLOWER CAR COVERS 2.667 MILES.